How to solve for x
-3/8(16x-24)=negative 9

Answer by stanbon(57361)

How to solve for x
-3/8(16x-24)=negative 9
-------------------------
(3/8)(16x-24) = 9
16x-24 = (8/3)9
16x-24 = 24
16x = 48
x = 3
==============
Cheers,
Stan H.
==============